# Quasi-boolean groups

### v2

@@ 6. $(vi)$
Every Boolean ring is an $\mathbb{F}_{2}$-algebra, so if $R$ is finite, then it is finitely generated as an $\mathbb{F}_{2}$-algebra, and hence $(i)$ implies $(ii)$. Every finitely generated $\mathbb{F}_{2}$-algebra is Noetherian by Hilbert’s basis theorem, so $(ii)$ implies $(iii)$. Every Boolean ring is zero dimensional by part $(ii)$ of Proposition 2.3, so if it is Noetherian, it is Artinian by a standard theorem in commutative algebra (see [1, Theorem 8.5 on page 90]). Therefore $(iii)$ implies $(iv)$. Every Artinian ring is a finite direct product of Artinian local rings (see [1, Theorem 8.7 on page 90]), so its spectrum is finite. Therefore $(iv)$ implies $(v)$. Now let $R$ be a Boolean ring whose spectrum is finite. Since $\mathrm{Spec}(R)$ is totally separated by Proposition 2.4, it is discrete, and hence $R\cong\mathbb{F}_{2}^{\mathrm{Spec}(R)}$ by Theorem 2.6. In particular, $(v)$ implies $(vi)$. If $R\cong\mathbb{F}_{2}^{X}$ for a finite set $X$, then $R$ is clearly finite, so $(vi)$ implies $(i)$. ∎
Notation 2.12.
@@
Let $R$ be a Boolean ring. We say that two elements $x,y\in R$ are orthogonal if $xy=0$. We say that an element $a\in R$ is an atom if it is non-zero and cannot be written as the sum of two non-zero orthogonal elements of $R$. The support of an element $x\in R$ is the subset $D(f)\subseteq\mathrm{Spec}(X)$ introduced in (2.4.1).
@@ 2. $(ii)$
    A non-zero element $a\in R$ is an atom if and only if its support cannot be written as the disjoint union of two non-empty open and closed subsets.
@@ 3. $(iii)$
    Every pair of different atoms of $R$ are orthogonal to each other.
@@
Note that the support of the product $xy$ is the intersection of the supports of $x$ and $y$. Since the support of an element of $R$ is empty if and only if it is zero by Theorem 2.6, claim $(i)$ follows. If $a=x+y$ such that $x,y$ are orthogonal and both non-zero, then the support of $a$ is the disjoint union of the supports of $x$ and $y$ by part $(i)$. On the other hand if the support of $a$ is the disjoint union of the non-empty open and closed subsets $X$ and $Y$, there are elements $x,y\in R$ whose support is $X$, $Y$, respectively, by Theorem 2.6. By part $(i)$ these are orthogonal, non-zero and their sum has the same support as $a$. So $a=x+z$, and hence claim $(ii)$ is true.
Let $a,b\in R$ be two atoms whose product is non-zero. Then $a=ab+a(1-b)$ and $aba(1-b)=a^{2}(b-b^{2})=0$, so $a(1-b)=0$ by the definition of atoms. We get that $a=ab$. The same reasoning for $b$ show that $b=ab$. Therefore $a=b$, and hence $(iii)$ holds. ∎
@@
Let $R$ be a finite Boolean ring. Then the atoms of $R$ form a natural basis of $R$ whose elements are orthogonal to each other. Moreover, every orthogonal basis consists of atoms.
@@
Now let $e_{1},e_{2},\ldots,e_{n}$ be an orthogonal basis. We need to show that each $e_{i}$ is an atom. We may assume without the loss of generality that $i=1$. Let $x$ be characteristic function of an element of the support of $e_{1}$. Then $x$ is an atom. The support of $e_{1}$ and $e_{i}$, $i\neq 1$, is disjoint since they are orthogonal. Hence $x$ and $e_{i}$ are orthogonal, so $xe_{i}=0$. Now write $x$ as a linear combination $\sum_{j\in J}e_{j}$ of the $e_{i}$. Since $0=xe_{i}=\sum_{j\in J}e_{j}e_{i}$ only if $i$ is not in $J$, the above argument implies $x=0$ or $x=e_{1}$. The former is not possible, since $x$ is not zero. Hence we get $x=e_{1}$, and $e_{1}$ is an atom. ∎
@@ 3. All about profinite spaces
We say that a topological space is profinite if it is the projective limit of discrete, finite topological spaces. Recall that a topological space $X$ is totally disconnected if for every point $x,y\in X$ the connected component of $x$ in $X$ is $x$ itself.
@@
First assume that $X$ satisfies $(i)$. Let $\mathcal{C}$ be a small category of discrete, finite topological spaces whose projective limit is $X$. Let $\mathrm{Ob}(\mathcal{C})$ denote the set of its objects, and let $\mathrm{Hom}_{\mathcal{C}}(A,B)$ denote the set of its morphisms for every $A,B\in\mathrm{Ob}(\mathcal{C})$. By definition, $X$ is homeomorphic to the closed subspace
@@
Next assume that $X$ satisfies $(ii)$. Since finite topological spaces are compact, their direct product is also compact by Tychonoff’s theorem. Since $X$ is a closed subspace of a compact space, it is compact, too. Moreover, the direct product of totally disconnected and Hausdorff topological spaces is totally disconnected and Hausdorff topological spaces. Since discrete topological spaces are totally disconnected and Hausdorff, and every subspace of a totally disconnected and Hausdorff topological space is also totally disconnected and Hausdorff, we get that the same holds for $X$, too. Therefore $X$ satisfies $(iii)$.
@@
Now consider the general case. By the above for every $x\in C$ there are disjoint open sets $U_{x}\subset X$ containing $x$ and $V_{x}\subset X$ containing $D$. Because $C$ is closed, it is compact, so there is a finite subset $x_{1},\ldots,x_{n}\in C$ such that $U_{x_{1}},\ldots,U_{x_{n}}$ cover $C$. Then $U=U_{x_{1}}\cup\cdots\cup U_{x_{n}}$ and $V=V_{x_{1}}\cap\cdots\cap V_{y_{n}}$ are disjoint open subsets containing $C$ and $D$, respectively. ∎
For every $x\in X$, let $Z(x)$ denote the set of all open and closed subsets of $X$ containing $x$, and let $Z_{x}$ denote the intersection of all elements of $Z(x)$. We need to show that if $y\in X$ is distinct from $x$ then $y\not\in Z_{x}$. Assume that this is not the case for some $y$. Then $Z_{x}$ contains at least two points, so it is not connected, since $X$ is totally disconnected. Therefore $Z_{x}$ is the disjoint union of two non-empty subsets $C,D\subset Z_{x}$ which are closed in $Z_{x}$.
Since $Z_{x}$ is the intersection of closed subsets, it is closed in $X$. Therefore $C$ and $D$ are also closed in $X$. Hence by Lemma 3.3 there exist disjoint open sets $U\subset X$ containing $C$ and $V\subset X$ containing $D$. Let $E$ be the complement of the union of $U$ and $V$ in $X$. It is closed, so it is compact since $X$ is Hausdorff. Since $U\cup V$ contains $Z_{x}$, the set $E$ is covered by the union of the complements of elements of $Z(x)$. Since $E$ is compact, it is already covered by the union of finitely many such sets. Therefore there is a finite subset $Z_{1},\ldots,Z_{n}\in Z(x)$ such that $Z=Z_{1}\cap\cdots\cap Z_{n}\subseteq U\cup V$.
Note that since each $Z_{i}$ is both open and closed, the same also holds for the finite intersection $Z$. Since each each $Z_{i}$ contains $x$, the set $Z$ also contains $x$. Both $Z\cap U$ and $Z\cap V$ are open in $Z$, and the latter is open, so $Z\cap U$ and $Z\cap V$ are open in $X$, too. The complement of $Z$ is open, so both $Z\cap U$ and $Z\cap V$ are the complement of the union of two open sets, so they are closed, too.
@@
For every positive integer $n\in\mathbb{N}$, let $\underline{n}$ denote the set $\{1,2,\ldots,n\}$ equipped with the discrete topology. Now let $X$ be an arbitrary topological space, and let $\mathcal{C}(X)$ denote the small category whose objects are continuous maps $f\colon X\to\underline{n}$ for some $n\in\mathbb{N}$, and the morphisms from an object $f\colon X\to\underline{n}$ to another object $g\colon X\to\underline{m}$ is an automatically continuous map $h\colon\underline{n}\to\underline{m}$ such that $g=h\circ f$.
@@
The profinite completion of $X$ is the projective limit $\widehat{X}$ of the small category $\mathcal{C}(X)$. It is equipped with a continuous map $u_{X}\colon X\to\widehat{X}$ satisfying the following universal property:
  1. (i)
    $\widehat{X}$ is profinite.
  2. (ii)
    For every continuous map $f\colon X\to Y$, where $Y$ is a discrete, finite topological space, there is a unique continuous map $\widehat{f}\colon\widehat{X}\to Y$ such that $f=\widehat{f}\circ u_{X}$.
This is because every discrete, finite topological space is homeomorphic to $\underline{n}$ for some $n\in\mathbb{N}$. Moreover $\widehat{X}$ is determined up to a unique homeomorphism by this universal property by the classical abstract non-sense argument.
It will be sufficient to show that $u_{X}$ is a homeomorphism when $X$ is compact and totally separated. Since $X$ is compact and $\widehat{X}$ is Hausdorff by the above, it will be sufficient to show that $u_{X}$ is a bijection. Let $x,y\in X$ be two different points. Since $X$ is totally separated, there is a continuous function $f\colon X\to\underline{2}$ such that $f(x)\neq f(y)$ as we saw in the proof of Theorem 2.8. Then $\widehat{f}(u_{X}(x))=f(x)\neq f(y)=\widehat{f}(u_{X}(y))$, so $u_{X}(x)\neq u_{X}(y)$. Therefore $u_{X}$ is injective. In order to see that $u_{X}$ is also surjective, we will need the following
@@
Note that it is enough to show the claim when $n=2$ as the general case follows by induction. Indeed let $n>2$ be such that we already know the claim for every $m<n$. Apply the $m=2$ case to the closed subsets $D_{1}=C_{1}$ and $D_{2}=C_{2}\cup\cdots\cup C_{n}$ to get two disjoint open and closed subsets $W_{1},W_{2}\subset X$ such that $D_{1}\subseteq W_{1}$, $D_{2}\subseteq W_{2}$, and $W_{1}\cup W_{2}=X$. Then apply the $m=n-1$ to $C_{2},\ldots,C_{n}$ inside $W_{2}$, which is possible since $W_{2}$ is closed in $X$, so it is compact and totally separated. Therefore there exist pairwise disjoint open and closed subsets $V_{2},\ldots,V_{n}\subset W_{2}$ such that $V_{i}$ contains $C_{i}$ for each $i\geq 2$, and $\cup_{i}V_{i}=W_{2}$. Since each $V_{i}$ is open and closed in an open and closed subset of $X$, it is also open and closed in $W_{2}$. Therefore $U_{1}=W_{1}$ and $U_{i}=V_{i}$ for $i\geq 2$ have the required properties.
@@
Now consider the $n=2$ case. By Lemma 3.3 there exist disjoint open sets $V_{1}\subset X$ containing $C_{1}$ and $V_{2}\subset X$ containing $C_{2}$. Because $X$ is totally separated, its open and closed sets form a subbasis for its topology, so for every $x\in C_{1}$ there is a closed and open subset $V_{x}\subset V_{1}$ containing $x$. Because $C_{1}$ is closed, it is compact, so there is a finite subset $x_{1},\ldots,x_{n}\in C$ such that $V_{x_{1}},\ldots,V_{x_{n}}$ cover $C$. Their union $U_{1}=V_{x_{1}}\cup\cdots\cup V_{x_{n}}$ is the union of open and closed subsets, so it is also open and closed, and it is contained in $V_{1}$. Its complement $U_{2}$ is also open and closed and contains $C_{2}$, so $U_{1}$ and $U_{2}$ have the required properties. ∎
@@
Now assume that $u_{X}$ is not surjective, so there is an $x\in\widehat{X}$ such that $x\not\in u_{X}(X)$. Since $X$ is compact and $\widehat{X}$ is Hausdorff by the above the image $u_{X}(X)$ is closed in $\widehat{X}$. So by Lemma 3.6 and since $X$ is compact and totally separated by the above, there are disjoint open and closed subsets $U_{1},U_{2}\subset\widehat{X}$ such that $u_{X}(X)\subseteq U_{1}$ and $x\in U_{2}$. Let $f:\widehat{X}\to\mathbb{F}_{2}$ the characteristic function of $U_{1}$. It is continuous, since $U_{1}$ is open and closed. Let $g:\widehat{X}\to\mathbb{F}_{2}$ be the identically $1$ function; it is also continuous, and $f\circ u_{X}=g\circ u_{X}$. However $f\neq g$, as $f(x)\neq g(x)$, which violates the universal property. This is a contradiction, so $(iv)$ implies $(i)$. ∎
Notation 3.7.
@@
Let $X$ be again an arbitrary topological space, and let $\beta\colon X\to\mathrm{Spec}(\mathbf{B}(X))$ be the map introduced in Definition 2.7. By Proposition 2.4 the space $\mathrm{Spec}(\mathbf{B}(X))$ is compact and totally separated, so it is profinite by Theorem 3.2. So it is a projective limit of finite, discrete spaces, and hence by the universal property of the profinite completion there is a unique continuous map $\widehat{\beta}\colon\widehat{X}\to\mathrm{Spec}(\mathbf{B}(X))$ such that $\beta=\widehat{\beta}\circ u_{X}$.
@@
According to the universal property of the profinite completion it will be sufficient to show that for every continuous map $f\colon X\to Y$, where $Y$ is a discrete, finite topological space, there is a unique continuous map $\widetilde{f}\colon\mathrm{Spec}(\mathbf{B}(X))\to Y$ such that $f=\widetilde{f}\circ\beta$. For every continuous map $m\colon T\to Q$ of topological maps let $m^{*}\colon\mathbf{B}(Q)\to\mathbf{B}(T)$ denote the induced ring homomorphism. Since $\beta^{*}\colon\mathbf{B}(X)\to\mathbf{B}(\mathrm{Spec}(\mathbf{B}(X))$ is an isomorphism there is a unique ring homomorphism $r\colon\mathbf{B}(Y)\to\mathbf{B}(\mathrm{Spec}(\mathbf{B}(X))$ such that $\beta^{*}\circ r=f^{*}$. By Theorem 2.10 there is a unique continuous map $\widetilde{f}\colon\mathrm{Spec}(\mathbf{B}(X))\to Y$ such that $r=\widetilde{f}^{*}$. Then $f^{*}=\beta^{*}\circ\widetilde{f}^{*}=(\widetilde{f}\circ\beta)^{*}$, so it will be sufficient to show that every continuous map $h\colon X\to Y$ is uniquely determined by $h^{*}$. Since for every $x\in X$ the point $h(x)$ is uniquely determined by the ideal

### v3

@@ 6. $(vi)$
Every Boolean ring is an $\mathbb{F}_{2}$-algebra, so if $R$ is finite, then it is finitely generated as an $\mathbb{F}_{2}$-algebra, and hence $(i)$ implies $(ii)$. Every finitely generated $\mathbb{F}_{2}$-algebra is Noetherian by Hilbert’s basis theorem, so $(ii)$ implies $(iii)$. Every Boolean ring is zero dimensional by part $(ii)$ of Proposition 2.3, so if it is Noetherian, it is Artinian by a standard theorem in commutative algebra (see [1, Theorem 8.5 on page 90]). Therefore $(iii)$ implies $(iv)$. Every Artinian ring is a finite direct product of Artinian local rings (see [1, Theorem 8.7 on page 90]), so its spectrum is finite. Therefore $(iv)$ implies $(v)$. Now let $R$ be a Boolean ring whose spectrum is finite. Since $\mathrm{Spec}(R)$ is totally separated by Proposition 2.5, it is discrete. This implies that $\mathbb{F}_{2}^{\mathrm{Spec}(R)}$ is isomorphic to $\mathbf{B}(\mathrm{Spec}(R)$ by the definition of the $\mathbf{B}(X)$ made in Example 2.2 since $\mathrm{Spec}(R)$ is discrete. By Theorem 2.7, $R$ is isomorphic to $\mathbf{B}(\mathrm{Spec}(R))$ via $\sigma$. Thus, we get $R\cong\mathbb{F}_{2}^{\mathrm{Spec}(R)}$. In particular, $(v)$ implies $(vi)$. If $R\cong\mathbb{F}_{2}^{X}$ for a finite set $X$, then $R$ is clearly finite, so $(vi)$ implies $(i)$. ∎
Notation 2.13.
@@
Let $R$ be a Boolean ring. We say that two elements $x,y\in R$ are orthogonal if $xy=0$. We say that an element $a\in R$ is an atom if it is non-zero and cannot be written as the sum of two non-zero orthogonal elements of $R$. The support of an element $x\in R$ is the subset $D(x)\subseteq\mathrm{Spec}(R)$ introduced in (2.5.1).
@@ 2. $(ii)$
    A non-zero element $a\in R$ is an atom if and only if its support cannot be written as the disjoint union of two non-empty open and closed subsets of $\mathrm{Spec}(R)$.
@@ 3. $(iii)$
    Every two distinct atoms of $R$ are orthogonal to each other.
@@
For $\mathfrak{p}\in\mathrm{Spec}(R)$, we have $xy\notin\mathfrak{p}$ if and only if $x\notin\mathfrak{p}$ and $y\notin\mathfrak{p}$. Hence $D(xy)=D(x)\cap D(y)$. Thus, the support of the product $xy$ is the intersection of the supports of $x$ and $y$. As explained in the first paragraph of the proof of Theorem 2.7, the nilradical of $R$ is zero and hence the intersection of all prime ideals of $R$ is zero by [1, Proposition 1.8 on page 5]. Hence, for $a\in R$, $D(a)=\emptyset$ if and only if $a=0$. This shows that the support of an element of $R$ is empty if and only if the element is zero. Claim $(i)$ now follows. By part $(iii)$ of Proposition 2.3, $(x,y)=(x+y)$, since $xy=0$. Hence, as in the proof of Theorem 2.7, $D(a)=D(x)\cup D(y)$. By part $(i)$, $D(x)\cap D(y)=\emptyset$. Therefore, $D(a)$ is the disjoint union $D(a)=D(x)\sqcup D(y)$. Also, as in the end of the proof of Proposition 2.5, $\mathrm{Spec}(R)=D(x)\sqcup D(1-x)$ and $\mathrm{Spec}(R)=D(y)\sqcup D(1-y)$, so $D(x)$ and $D(y)$ are open and closed subsets of $\mathrm{Spec}(R)$. Now assume that the support of $a$ is the disjoint union of the non-empty open and closed subsets $X$ and $Y$ of $\mathrm{Spec}(R)$. Let $f_{X},f_{Y}\colon\mathrm{Spec}(R)\to\mathbb{F}_{2}$ be the characteristic functions of $X$ and $Y$, respectively, which are continuous by the assumption that $X$ and $Y$ are open and closed subsets of $\mathrm{Spec}(R)$. Then, as in the proof of Theorem 2.7, there exist $x,y\in R$ such that $X=D(f_{X})=D(x)$ and $Y=D(f_{Y})=D(y)$. Since $D(\sigma(a))=D(a)=D(x)\sqcup D(y)=D(x+y)=D(\sigma(x+y))$, we have $\sigma(a)=\sigma(x+y)$, so by Theorem 2.7, $a=x+y$. Let $a,b\in R$ be two atoms whose product is non-zero. Then a=ab+a(1b)a=ab+a(1-b) and $aba(1-b)=a^{2}(b-b^{2})=0$, so a(1b)=0a(1-b)=0 by the definition of atoms. We get that $a=ab$. The same reasoning for $b$ shows that $b=ab$. Therefore $a=b$, and hence $(iii)$ holds. ∎
@@
Let $R$ be a finite Boolean ring. Then the atoms of $R$ form a natural basis of $R$ as an $\mathbb{F}_{2}$-vector space whose elements are orthogonal to each other. Moreover, every orthogonal basis consists of atoms.
@@
Now let $e_{1},e_{2},\ldots,e_{n}$ be an orthogonal basis. We need to show that each $e_{i}$ is an atom. We may assume without the loss of generality that $i=1$. Let $x\in R$ be such that $\sigma(x)$ is the characteristic function of an element of the support of $e_{1}$. Then $x$ is an atom. The supports of $e_{1}$ and $e_{i}$, $i\neq 1$, are disjoint since they are orthogonal by claim $(i)$ of Lemma 2.16. Hence $x$ and $e_{i}$ are orthogonal, so $xe_{i}=0$. Now write $x$ as a linear combination $\sum_{j\in J}e_{j}$ of the $e_{i}$ for $J\subseteq\{1,\ldots,n\}$. Since $0=xe_{i}=\sum_{j\in J}e_{j}e_{i}$ only if $i$ is not in $J$, the above argument implies $x=0$ or $x=e_{1}$. The former is not possible, since $x$ is not zero. Hence we get $x=e_{1}$, and $e_{1}$ is an atom by the first paragraph of the proof. ∎
@@ 3. All about profinite spaces
We say that a topological space is profinite if it is the projective limit of discrete, finite topological spaces. Recall that a topological space $X$ is totally disconnected if for every point $x\in X$ the connected component of $x$ in $X$ is $\{x\}$.
@@
First assume that $X$ satisfies $(i)$. Recall that a category is called small if both its class of objects and class of morphisms are sets and not proper classes. Let $\mathcal{C}$ be a small category of discrete, finite topological spaces whose projective limit is $X$. Let $\mathrm{Ob}(\mathcal{C})$ denote the set of its objects, and let $\mathrm{Hom}_{\mathcal{C}}(A,B)$ denote the set of its morphisms for every $A,B\in\mathrm{Ob}(\mathcal{C})$. By definition, $X$ is homeomorphic to the closed subspace
@@
Next assume that $X$ satisfies $(ii)$. Since finite topological spaces are compact, their direct product is also compact by Tychonoff’s theorem. Since $X$ is a closed subspace of a compact space, it is compact, too. Moreover, the direct product of totally disconnected and Hausdorff topological spaces is totally disconnected and Hausdorff. Since discrete topological spaces are totally disconnected and Hausdorff, and every subspace of a totally disconnected and Hausdorff topological space is also totally disconnected and Hausdorff, we get that the same holds for $X$, too. Therefore $X$ satisfies $(iii)$.
@@
Now consider the general case. By the above for every $x\in C$ there are disjoint open sets $U_{x}\subset X$ containing $x$ and $V_{x}\subset X$ containing $D$. Because $C$ is closed, it is compact, so there is a finite subset $\{x_{1},\ldots,x_{n}\}\subset C$ such that $U_{x_{1}},\ldots,U_{x_{n}}$ cover $C$. Then $U\coloneqq U_{x_{1}}\cup\cdots\cup U_{x_{n}}$ and $V\coloneqq V_{x_{1}}\cap\cdots\cap V_{x_{n}}$ are disjoint open subsets containing $C$ and $D$, respectively. ∎
For every $x\in X$, let $Z(x)$ denote the set of all open and closed subsets of $X$ containing $x$, and let $Z_{x}$ denote the intersection of all elements of $Z(x)$. By the definition of a totally separated topological space, we need to show that if $y\in X$ is distinct from $x$ then $y\not\in Z_{x}$. Assume that this is not the case for some $y$. Then $Z_{x}$ contains at least two points, so it is not connected, since $X$ is totally disconnected. Therefore $Z_{x}$ is the disjoint union of two non-empty subsets $C,D\subset Z_{x}$ which are closed in $Z_{x}$. Since $Z_{x}$ is the intersection of closed subsets, it is closed in $X$. Therefore $C$ and $D$ are also closed in $X$. Hence by Lemma 3.3, we can find an open set $U\subset X$ containing $C$ such that $\overline{U}\cap D=\emptyset$. Let $E$ be the complement of the union of $U$ and $V$ in $X$. It is closed in $X$, so it is compact since $X$ is compact. Since $U\cup V$ contains $Z_{x}$, the set $E$ is covered by the union of the complements of elements of $Z(x)$. Since $E$ is compact, it is already covered by the union of finitely many such sets. Therefore there are finitely many sets $Z_{1},\ldots,Z_{n}\in Z(x)$ such that $Z=Z_{1}\cap\cdots\cap Z_{n}\subseteq U\cup V$.
Note that since each $Z_{i}$ is both open and closed, the same also holds for the finite intersection $Z$. Since each $Z_{i}$ contains $x$, the set $Z$ also contains $x$. Both $Z\cap U$ and $Z\cap V$ are open in $Z$, and $Z$ is open in $X$, so $Z\cap U$ and $Z\cap V$ are open in $X$, too. Since $Z$ lies in the disjoint union of $U$ and $V$, every element of $X$ which is not in $Z\cap U$ is either not in $Z$ or it is in $Z\cap V$. Thus, we get $Z\cap U=X-[(X-Z)\cup(Z\cap V)]$. Since $Z$ is closed, XZX-Z is open and we just showed $Z\cap V$ is open. Hence $Z\cap U$ is also closed since it is the complement of an open subset. The same argument with the roles of $U$ and $V$ reversed shows that also $Z\cap V$ is closed.
@@
Let $X$ be a topological space. We denote by $\mathcal{R}(X)$ the set of open and closed equivalence relations $R$ on $X$ such that $X/R$ is finite and discrete. We may view $R\in\mathcal{R}(X)$ as an open and closed subset of $X\times X$ and we order elements in $\mathcal{R}(X)$ by inclusion, i.e., $R_{1}\geq R_{2}$ in $\mathcal{R}(X)$ if and only if $R_{1}\subseteq R_{2}$. This turns $\mathcal{R}(X)$ into a directed partially ordered set. To see that $\mathcal{R}(X)$ is directed, let $R_{1},R_{2}\in\mathcal{R}(X)$. Then $R_{1}\cap R_{2}\subset X\times X$ is open and closed. Moreover, the quotient maps $X/(R_{1}\cap R_{2})\to X/R_{1}$ and $X/(R_{1}\cap R_{2})\to X/R_{2}$ are continuous and induce a continuous map $X/(R_{1}\cap R_{2})\to X/R_{1}\times X/R_{2}$. This map is injective since $(x,y)\in R_{1}$ and $(x,y)\in R_{2}$ if and only if $(x,y)\in R_{1}\cap R_{2}$. Thus, we can consider $X/(R_{1}\cap R_{2})$ as a subspace of the finite and discrete space $X/R_{1}\times X/R_{2}$, and hence $X/(R_{1}\cap R_{2})$ is finite and discrete, too. This shows $R_{1}\leq R_{1}\cap R_{2}$ and $R_{2}\leq R_{1}\cap R_{2}$ as required.
@@
Let $X$ be a topological space. We define the profinite completion of $X$ to be the profinite space $\widehat{X}$ given by the projective limit $\widehat{X}\coloneqq\lim_{R\in\mathcal{R}(X)}X/R$ in the category of topological spaces. It is equipped with a continuous map $u_{X}\colon X\to\widehat{X}$ induced by the quotient maps $X\to X/R$ and the general properties of projective limits.
Remark 3.6.
The continuous map $u_{X}\colon X\to\widehat{X}$ satisfies the following universal property: For every continuous map $f\colon X\to Y$, where $Y$ is a discrete and finite topological space, there is a unique continuous map $\widehat{f}\colon\widehat{X}\to Y$ such that $f=\widehat{f}\circ u_{X}$. This follows from the fact that $f$ induces an equivalence relation $R_{f}\subset X\times X$ on $X$ given by $(x,x^{\prime})\in R_{f}$ if and only if $f(x)=f(x^{\prime})$. For $x\in X$, let $y=f(x)$. Since $Y$ is discrete, the preimage $f^{-1}(y)$ is both open and closed in $X$. Moreover, since $Y$ is finite, there are finitely many equivalence classes under $R_{f}$. Hence the quotient space $X/R_{f}$ is finite and discrete. This shows that $R_{f}\in\mathcal{R}(X)$. Let $\overline{f}\colon X/R_{f}\to Y$ be the induced map. The map $\widehat{f}$ is now defined as the composition of the canonical map $\widehat{X}\to X/R_{f}$ and the map $\overline{f}$.
Now we return to the proof of Theorem 3.2. To show that $(iv)$ implies $(i)$ it will be sufficient to show that $u_{X}$ is a homeomorphism when $X$ is compact and totally separated. Since $X$ is compact and $\widehat{X}$ is Hausdorff by $(i)\Rightarrow(iii)$ of Theorem 3.2 which we have already shown, it will be sufficient to show that $u_{X}$ is a bijection by [2, I §9.4, Corollary 2 on page 87] which states that a continuous bijection from a compact space to a Hausdorff space is a homeomorphism. Let $x,y\in X$ be two distinct points. Since $X$ is totally separated, there is a continuous function $f\colon X\to\mathbb{F}_{2}$ such that $f(x)\neq f(y)$ as we saw in the proof of Theorem 2.9. Then $\widehat{f}(u_{X}(x))=f(x)\neq f(y)=\widehat{f}(u_{X}(y))$, so $u_{X}(x)\neq u_{X}(y)$. Therefore $u_{X}$ is injective. In order to see that $u_{X}$ is also surjective, we will need the following
@@
We first show that it is enough to prove the claim when $n=2$ as the general case follows by induction. Indeed let $n>2$ be such that we already know the claim for every $m<n$. Apply the case $m=2$ to the closed subsets $D_{1}=C_{1}$ and $D_{2}=C_{2}\cup\cdots\cup C_{n}$ to get two disjoint open and closed subsets $W_{1},W_{2}\subset X$ such that $D_{1}\subseteq W_{1}$, $D_{2}\subseteq W_{2}$, and $W_{1}\cup W_{2}=X$. Then apply the case m=n1m=n-1 to $C_{2},\ldots,C_{n}$ inside $W_{2}$, which is possible since $W_{2}$ is closed in $X$, so it is compact and totally separated. Therefore there exist pairwise disjoint open and closed subsets $V_{2},\ldots,V_{n}\subset W_{2}$ such that $V_{i}$ contains $C_{i}$ for each $i\geq 2$, and $\cup_{i}V_{i}=W_{2}$. Since each $V_{i}$ is open and closed in an open and closed subset of $X$, it is also open and closed in $X$. Therefore $U_{1}=W_{1}$ and $U_{i}=V_{i}$ for $i\geq 2$ have the required properties.
@@
Now consider the case $n=2$. Since $X$ is totally separated by assumption, it is also Hausdorff. By Lemma 3.3, there then exist disjoint open sets $V_{1}\subset X$ containing $C_{1}$ and $V_{2}\subset X$ containing $C_{2}$. Because $X$ is totally separated, its open and closed sets form a subbasis for its topology, so for every $x\in C_{1}$ there is a closed and open subset $V_{x}\subset V_{1}$ containing $x$. Because $C_{1}$ is closed, it is compact, so there is a finite subset $x_{1},\ldots,x_{n}\in C$ such that $V_{x_{1}},\ldots,V_{x_{n}}$ cover $C$. Their union $U_{1}=V_{x_{1}}\cup\cdots\cup V_{x_{n}}$ is the union of open and closed subsets, so it is also open and closed, and it is contained in $V_{1}$. Its complement $U_{2}$ is also open and closed and contains $C_{2}$, so $U_{1}$ and $U_{2}$ have the required properties. ∎
@@
Now assume that $u_{X}$ is not surjective, so there is an $x\in\widehat{X}$ such that $x\not\in u_{X}(X)$. Since $X$ is compact and $\widehat{X}$ is Hausdorff by $(i)\Rightarrow(iii)$ of Theorem 3.2 which we have already proven, the image $u_{X}(X)$ is closed in $\widehat{X}$ by [2, I §9.4, Corollary 2 on page 87] which states that a continuous map from a compact space to a Hausdorff space is closed. So by Lemma 3.7 and since $\widehat{X}$ is compact and totally separated by $(i)\Rightarrow(iv)$ of Theorem 3.2, which we already have proved, there are disjoint open and closed subsets $U_{1},U_{2}\subset\widehat{X}$ such that $u_{X}(X)\subseteq U_{1}$ and $x\in U_{2}$. Let $f:\widehat{X}\to\mathbb{F}_{2}$ the characteristic function of $U_{1}$. It is continuous, since $U_{1}$ is open and closed. Let $g\colon\widehat{X}\to\mathbb{F}_{2}$ be the identically $1$ function. It is continuous, and $f\circ u_{X}=g\circ u_{X}$. However $f\neq g$, as $f(x)\neq g(x)$, which violates the universal property. This is a contradiction, so $(iv)$ implies $(i)$. ∎
Remark 3.8.
We note that the profinite completion of $X$ as a topological space in Definition 3.5 differs from the profinite completion of $X$ as a set. The latter is the completion of $X$ as a space with the discrete topology. For example, if $X$ is connected then the only non-empty open and closed equivalence relation in $\mathcal{R}(X)$ is $X\times X$. As a set, however, there may be many equivalence relations on $X$ such that $X/R$ is finite. Moreover, if $X$ is a profinite space, the proof of $(iv)\Rightarrow(i)$ of Theorem 3.2 shows that $u_{X}$ is a homeomorphism. This may not be the case if we first equip $X$ with the discrete topology.
Next, we show that profinite completion and the composition of functors $\mathrm{Spec}\circ\mathbf{B}$ of Section 2 yield homeomorphic spaces.
Notation 3.9.
@@
Let $X$ be again an arbitrary topological space, and let $\beta\colon X\to\mathrm{Spec}(\mathbf{B}(X))$ be the map introduced in Notation 2.8. By Proposition 2.5 the space $\mathrm{Spec}(\mathbf{B}(X))$ is compact and totally separated, so it is profinite by $(iv)\Rightarrow(i)$ of Theorem 3.2. So it is a projective limit of finite, discrete spaces, and hence by the universal property of the profinite completion there is a unique continuous map $\widehat{\beta}\colon\widehat{X}\to\mathrm{Spec}(\mathbf{B}(X))$ such that $\beta=\widehat{\beta}\circ u_{X}$.
@@
According to the universal property of the profinite completion it will be sufficient to show that for every continuous map $f\colon X\to Y$, where $Y$ is a discrete, finite topological space, there is a unique continuous map $\widetilde{f}\colon\mathrm{Spec}(\mathbf{B}(X))\to Y$ such that $f=\widetilde{f}\circ\beta$. For every continuous map $m\colon T\to Q$ of topological spaces, let $m^{*}\colon\mathbf{B}(Q)\to\mathbf{B}(T)$ denote the induced ring homomorphism. Since $\beta^{*}\colon\mathbf{B}(\mathrm{Spec}(\mathbf{B}(X)))\to\mathbf{B}(X)$ is the inverse of $\sigma$ in Notation 2.6 for $R=\mathbf{B}(X)$, $\beta^{*}$ is an isomorphism by Theorem 2.7. Hence there is a unique ring homomorphism $r\colon\mathbf{B}(Y)\to\mathbf{B}(\mathrm{Spec}(\mathbf{B}(X)))$ such that $\beta^{*}\circ r=f^{*}$. By Theorem 2.11 there is a unique continuous map $\widetilde{f}\colon\mathrm{Spec}(\mathbf{B}(X))\to Y$ such that $r=\widetilde{f}^{*}$. Then $f^{*}=\beta^{*}\circ\widetilde{f}^{*}=(\widetilde{f}\circ\beta)^{*}$, so it will be sufficient to show that every continuous map $h\colon X\to Y$ is uniquely determined by $h^{*}$. Since for every $x\in X$ the point $h(x)$ is uniquely determined by the ideal